# Cooperative Edge Caching Based on Elastic Federated and Multi-Agent Deep Reinforcement Learning in Next-Generation Networks

### v1

@@ I Introduction
In recent years, with the increasing popularity of smart devices, we have witnessed an unprecedented growth in mobile data traffic, which has imposed a heavy burden on wireless networks[1, 2]. As users increasingly rely on user equipment (UEs) like mobile devices and home routers to access content from wireless networks, it becomes challenging for ensuring a satisfactory quality of service to meet their demands. To address this challenge, edge caching has emerged as a promising solution for next-generation networks [3, 4]. Through the implementation of caching units in wireless edge nodes, such as small-cell base stations (SBSs), UEs can fetch users’ desired contents from nearby SBSs instead of remote servers or cloud. This process significantly reduces traffic load, alleviates network congestion, reduces latency and improves system performance[5].
@@
After each local SBS has predicted the popular contents, deploying edge caches poses an additional challenge. Current caching methods in wireless networks often worked separately. Each SBS decides where to cache contents based on certain criteria. While there have been efforts to study collaborative caching, there is still a lack of research in making caching decisions among SBSs based on popular contents. In next-generation networks, the importance of collaborative caching is further strengthened. Here, cached contents can be shared among adjacent SBSs via the Xn interface [11]. This allows UEs to indirectly obtain the requested contents from the adjacent SBSs even if they cannot directly fetch the requested contents from the local SBS. However, the number of predicted popular contents for each SBS is usually larger than the limited cache capacity of the SBS. Therefore, this requires cooperation between SBSs to cache popular contents. Additionally, the placement of these popular contents in the right SBSs is also critical, as it directly impacts the cost of fetching contents across the network. In addition, the dimensions of caching decision will increase with the number of SBSs and cached contents, inevitably adding complexity to the caching problem in the entire network. This makes it challenging to coordinate caching among SBSs to reduce the cost for fetching contents in the next-generation network. Multi-agent reinforcement learning (MADRL) is a suitable method that can overcome the above problem, and has begun to be used in cooperative edge caching. Instead of the traditional way of making separate decisions, MADRL promotes cooperation among SBSs. Each SBS, which is considered an agent, does not just decide where to cache popular contents based on its own prediction. Instead, it should work together with other SBSs, taking into account their decisions to create a network-wide caching strategy driven by content popularity. If some UEs’ demands change, MADRL can adjust the caching decisions. This adaptability is crucial for handling the increased complexity that comes with the growing caching needs in the next-generation network.
In this paper, we propose a cooperative edge caching scheme based on elastic federated learning and multi-agent deep reinforcement learning (CEFMR) scheme to optimize the cost of fetching contents in the next-generation network. The contributions of this paper are summarized as follows [^1] The source code has been released at: https://github.com/qiongwu86/Edge-Caching-Based-on-Multi-Agent-Deep-Reinforcement-Learning-and-Federated-Learning
@@
The remainder of this paper is structured as follows. Section II reviews the related works on contents edge caching based on the MADRL and FL algorithms. Section III provides a concise overview of the system model. Section IV introduces the proposed cooperative edge caching scheme. We present some simulation results in Section V, and then conclude them in Section VI.
@@ II Related Work
We first review some current existing works on content caching schemes based on the MADRL algorithm. In [13], Araf et al. employed a multi-agent actor-critic reinforcement learning approach, combined with UAV (Unmanned Aerial Vehicle) assistance, to tackle the problem of optimizing cooperative caching strategies on the network edge. In [14], Chen et al. proposed a multi-agent actor-critic framework with variational recurrent neural networks (VRNN) to estimate content popularity. Through multi-agent deep reinforcement learning, the cooperative edge caching algorithm substantially improved the benefits of edge caches in ultra-dense next-generation networks. In [15], Jiang et al. employed multi-agent reinforcement learning to design content caching strategies in mobile device-to-device (D2D) networks, addressing the challenge of content caching without prior knowledge of popularity distribution. The proposed approach, using Q-learning, maximized caching rewards by treating users as agents and caching decisions as actions. In [16], Wu et al. proposed the integration of cooperative coded caching using maximum distance separable (MDS) codes in small cell networks to alleviate traffic loads. To adapt to dynamic and unknown content popularity, a multi-agent deep reinforcement learning framework was introduced to intelligently update cached content, with a focus on minimizing long-term expected front-haul traffic loads, resulting in improved caching efficiency. In [17], Zhong et al. investigated cache placement by using the collaborative MADRL algorithm in ultra-dense small cell networks (SCNs) to maximize cache hit rates in centralized and decentralized scenario settings. In [18], Jiang et al. explored edge caching for the internet of vehicles (IoVs) using distributed MARL. They introduced a hierarchical caching architecture to minimize content delivery overhead, extended single-agent reinforcement learning to a multi-agent system, and achieved superior performance in comparison to other caching strategies through simulation results. In [21], Song et al. addressed the content caching and sharing problem in cooperative base stations while considering unknown content popularity distribution and content retrieval costs. This paper formulated the problem as a multi-armed bandit (MAB) learning problem, jointly optimizing caching and sharing strategies. Two algorithms, a centralized approach using semidefinite relaxation and a decentralized approach based on the alternating direction method of multipliers (ADMM), were proposed to solve the problem efficiently. In [22], Cai et al. proposed a cross-tier cooperative caching architecture for all content that allows distributed cache nodes to cooperate, using a MADRL approach to model the decision process between heterogeneous cache nodes. In [23], Wang et al. introduced an intelligent edge caching framework named MacoCache, which incorporated MADRL to enhance edge caching performance in distributed environments. By employing the advanced actor-critic method and integrating long short-term memory (LSTM) for time series dynamics, MacoCache successfully reduced latency by an average of 21% and costs by 26% compared to existing learning-based caching solutions. In [24], Chen et al. utilized a refined MADRL tailored for the 5G network structure to improve scalable video coding (SVC)-based caching and video quality service strategies. The introduced value-decomposed dimensional network (VDDN) algorithm and a dimension decomposition method in the dueling deep Q-network addressed the challenges of large action spaces and computational efficiency. In [25], Zhang et al. introduced an edge caching model for IoV content distribution, allowing vehicles to collaboratively cache content and account for varying content popularity and channel conditions. An autonomous decision-making approach, an edge caching approach for the IoV based on multi-agent deep reinforcement learning, was proposed where vehicles act as agents to optimize caching decisions. However, the above works predominantly relied on the direct application of MADRL for content caching challenges. It is worth noting that these works did not particularly emphasize the complex challenges related to content prediction, especially the prediction of content popularity and training with machine learning.
@@
Recently, there have been several works on designing content caching schemes based on MADRL and FL. In [26], Chang et al. introduced a federated multi-agent reinforcement Learning (FMRL) method for optimizing edge caching. Using a Markov decision process (MDP) and a specialized regression model, it predicted content popularity, improved server rewards, and optimized cache decision-making. An enhanced FL technique aggregated local models, improving training and preserving user privacy. In [27], Prathiba et al. introduced the federated Learning and edge cache-assisted Cybertwin framework in response to rising autonomous vehicle (AV) technology and mobile traffic. This framework integrated the cybertwin concept to merge physical and digital systems. It employed the federated MADRL algorithm for optimized learning, considering various factors. Additionally, the federated reinforcement learning-based edge caching algorithm was used for efficient caching in 6G vehicle-to-everything networks. In [28], Sun et al. employed FL to perceive user preferences across various districts without exposing user requests. A content blocks-based RL is utilized for intelligent caching, taking into account the receiving scenarios of AVs. By integrating FL with MARL, this paper achieved distributed training of local models, which not only reduced the state-action space dimension but also circumvented local optimal solutions. However, the above works did not consider the personalization of each UE in FL training and the popular contents for different SBSs.
To sum up, no work has considered the personalized model for each UE to design content caching schemes based on MADRL and FL, which motivated our research.
@@
$\bm{a^{t}}$
the global action at time slot $t$.
$a_{b}^{t}$
the action of SBS $b$ at time slot $t$.
$B$
the number of SBSs.
$b_{d}$
the bias vector of the decoder network.
$b_{e}$
the bias vector of the encoder network.
$C$
the cache capacity of each SBS.
$c_{b}^{t}$
@@
the current contents cached by SBS $b$ at time slot $t$.
$d$
the size of entire data of all UEs within the coverage of SBS.
$d(\cdot)$
@@
the non-linear and logically-activated function of the discriminator network.
$d_{i}$
the size of the training data of each UE.
$D$
the replay buffer.
$D(z)$
the confidence of $z$.
$e$
the number of iteration of the local AAE training for UE.
$E^{\prime}$
@@
the number of episodes in the testing phase of MADRL algorithm.
$F_{p}$
the number of predicted popular contents.
$J(\bm{\pi^{t}})$
@@
the expected long-term discounted reward at time slot $t$.
$|L|$
the number of layers of the AAE model.
$M$
the number of transition tuples in a mini-batch.
$n_{i,k}^{r}$
local training data set of UE $i$ in the iteration $k$ of round $k$.
$p(\cdot)$
@@
the non-linear and logically-activated function of the decoder network.
$P(z)$
the predefined distribution.
$p_{b}$
the predicted popular contents of SBS $b$ coverage.
$q(\cdot)$
@@
the non-linear and logically-activated function of the encoder network.
$R^{t}$
the global reward at the time slot $t$.
$R_{b}^{t}$
the local reward of SBS $b$ at time slot $t$.
$R_{max}$
@@
the number of rounds to train the AAE model in the elastic FL.
$r_{b}^{r}(t)$
@@
the number of contents that SBS $b$ fetches from the local cache at time slot $t$.
$r_{b,c}^{r}(t)$
the number of contents that SBS $b$ fetches from CS at time slot $t$.
$r_{b,e}^{r}(t)$
the contents replaced in SBS $b$ during time slot $t$.
$r_{b,n}^{r}(t)$
@@
the number of contents that SBS $b$ fetches from adjacent SBSs at time slot $t$.
$s_{b}^{t}$
the local state of SBS $b$ at time slot $t$.
$\bm{s^{t}}$
the global state at time slot $t$.
$T$
@@
the number of time slots of the MADRL algorithm.
$X_{i}^{r}$
the rating matrix of UE $i$ in round $r$.
$\tilde{X}_{i}^{r}$
the reconstructed rating matrix of UE $i$ in round $r$.
$\alpha$
@@
the cost of fetching the contents from local SBS $b$.
$\alpha_{i}$
the elastic parameter of UE $i$.
$\beta$
@@
the cost that the UE within the coverage of local SBS $b$ fetch the contents from adjacent SBSs.
$\chi$
the cost of fetching the contents from CS.
$\delta$
the cost of replacing a content.
$\gamma$
the discount factor.
$\eta$
the fixed learning rate.
$\theta_{b}$
the parameters of the actor network of agent $b$.
$\theta_{b}^{\prime}$
@@
the parameters of the target critic network of agent $b$.
$\varphi$
the parameters of the global critic network.
$\varphi^{\prime}$
the parameters of the target global critic network.
$\tau$
@@
the constant is used to update the target networks.
$\omega^{r}$
the global AAE model in round $r$.
$\omega_{e}^{r}$
the global encoder network in round $r$.
$\omega_{d}^{r}$
the global decoder network in round $r$.
$\omega_{d^{\prime}}^{r}$
@@
the global discriminator network in round $r$.
$\omega_{i}^{r}$
the local model of UE $i$ in round $r$.

### v2

@@ I Introduction
In recent years, with the increasing popularity of smart devices, we have witnessed an unprecedented growth in mobile data traffic, which has imposed a heavy burden on wireless networks[1, 2, 3]. As users increasingly rely on user equipment (UEs) like mobile devices and home routers to access content from wireless networks, it becomes challenging for ensuring a satisfactory quality of service to meet their demands[6, 5, 4]. To address this challenge, edge caching has emerged as a promising solution for next-generation networks [7, 8, 9]. Through the implementation of caching units in wireless edge nodes, such as small-cell base stations (SBSs), UEs can fetch users’ desired contents from nearby SBSs instead of remote servers or cloud. This process significantly reduces traffic load, alleviates network congestion, reduces latency and improves system performance[10].
@@
After each local SBS has predicted the popular contents, deploying edge caches poses an additional challenge. In next-generation networks, the importance of collaborative caching is further strengthened. Here, cached contents can be shared among adjacent SBSs via the Xn interface [16]. This allows UEs to indirectly obtain the requested contents from the adjacent SBSs even if they cannot directly fetch the requested contents from the local SBS. The majority of research in collaborative caching has focused on joint optimization with metrics like resource allocation and latency reduction, rather than concentrating on popular contents. Additionally, while studies employing artificial intelligence technology generally use popular contents in their caching strategies, they often overlook the crucial aspect of predicting content popularity for future proactive caching. Therefore, predicting future popular contents and utilizing them for collaborative caching can significantly enhance caching efficiency. However, the number of predicted popular contents for each SBS is usually larger than the limited cache capacity of the SBS. Therefore, this requires cooperation between SBSs to cache popular contents. Additionally, the placement of these popular contents in the right SBSs is also critical, as it directly impacts the cost of fetching contents across the network. In addition, the dimensions of caching decision will increase with the number of SBSs and cached contents, inevitably adding complexity to the caching problem in the entire network. This makes it challenging to coordinate caching among SBSs to reduce the cost for fetching contents in the next-generation network. Multi-agent reinforcement learning (MADRL) is a suitable method that can overcome the above problem, and has begun to be used in cooperative edge caching. Instead of the traditional way of making separate decisions, MADRL promotes cooperation among SBSs. Each SBS, which is considered an agent, does not just decide where to cache popular contents based on its own prediction. Instead, it should work together with other SBSs, taking into account their decisions to create a network-wide caching strategy driven by content popularity. If some UEs’ demands change, MADRL can adjust the caching decisions. This adaptability is crucial for handling the increased complexity that comes with the growing caching needs in the next-generation network.
In this paper, we propose a cooperative edge caching scheme based on elastic federated learning and multi-agent deep reinforcement learning (CEFMR) scheme to optimize the cost of fetching contents in the next-generation network. The contributions of this paper are summarized as follows [^1] This paper has been submitted to IEEE TNSM. The source code has been released at: https://github.com/qiongwu86/Edge-Caching-Based-on-Multi-Agent-Deep-Reinforcement-Learning-and-Federated-Learning
@@
The remainder of this paper is structured as follows. Section II reviews the related works on the research on popularity prediction and collaborative caching. Section III provides a concise overview of the system model. Section IV introduces the proposed cooperative edge caching scheme. We present some simulation results in Section V, and then conclude them in Section VI.
@@ II Related Work
We first reviewed the research on popularity prediction and collaborative caching, and then we reviewed the studies on collaborative caching based on MADRL and FL.
Currently, there are studies on content popularity prediction. In [17], Pervej et al. proposed a novel approach to edge caching in next-generation wireless networks using a long short-term memory based sequential model to dynamically predict user preferences for content caching. Additionally, they formulated a non-convex optimization problem to minimize content sharing costs and employed a greedy algorithm to find a sub-optimal solution. In [18], Yao et al. introduced vehicular content centric network to address mobility challenges in vehicular ad hoc networks by implementing content centric networking. They proposed popularity-based content caching, a cache replacement scheme that predicts content popularity using hidden Markov Model based on received interests, request ratio, frequency and content priority. In [19], Tao et al. proposed a Gaussian process based Poisson regressor to accurately and efficiently model content request patterns. Bayesian learning is employed to robustly learn model parameters, and to handle the lack of closed-form expression for the posterior distribution, a Stochastic Variance Reduced Gradient Hamiltonian Monte Carlo method is applied. In [20], Liu et al. presented a content popularity prediction algorithm based on auto-regressive models for Information-Centric Networking in the Internet of Things, improving cache hit rates, reducing network traffic, and minimizing service access delays, particularly beneficial for real-time streaming media services. In [21], Feng et al. introduced a content popularity prediction algorithm, GRU-Attention, tailored for Vehicle Named Data Networking, which relied on in-network caching to swiftly serve subsequent content requests. By leveraging the attention mechanism and GRU model, it accurately forecasted future popular contents based on multiple historical request patterns, enhancing prediction accuracy and optimizing cache utilization within the network.
There are already studies on collaborative caching. In [22], Pervej et al. tackled the optimization of collaborative caching in heterogeneous edge networks by developing an improved Particle Swarm Optimization algorithm, demonstrating through numerical analysis and simulations that this approach significantly enhances cache hit ratios compared to traditional caching strategies. In [23], Saputra et al. employed a cooperative mobile edge caching network by jointly optimizing collaborative caching and routing decisions to reduce content access delays and backhaul traffic. To address the optimization challenge, they transformed the problem into a mixed-integer nonlinear programming problem and utilized a branch-and-bound algorithm with the interior-point method to find a near-optimal solution. In [24], Xie et al. proposed a cooperative caching scheme where user terminals cache popular services to an intelligent routing relay by using an online reverse auction combined with a first-come-first-served strategy. In [25], Li et al. employed a crowdsourced approach for cooperative caching in Content-Centric Mobile Networking. They introduced a robust caching control scheme using Kullback-Leibler divergence to manage uncertainties in user mobility, finally formulated as a chance-constrained robust optimization problem. In [26], Liao et al. employed cooperative caching strategies using maximum distance separable codes for content restructuring and optimizes both content placement and cooperation policies among base stations. A compound caching strategy, namely multicast-aware cooperative caching assuming fixed and dynamic cooperative policies, respectively, was developed to combine the merits of multicast-aware content delivery and cooperative content sharing. In [27], Fang et al. investigated the economic interactions between Internet service providers and content providers in wireless networks, focusing on optimizing content delivery through edge caches and content popularity. They proposed a centralized model to solve the profit split problem and employed a Stackelberg game for a distributed solution.
@@
Furthermore, there are already studies on collaborative caching based on MADRL. In [28], Araf et al. employed a multi-agent actor-critic reinforcement learning approach, combined with UAV (Unmanned Aerial Vehicle) assistance, to tackle the problem of optimizing cooperative caching strategies on the network edge. In [29], Chen et al. proposed a multi-agent actor-critic framework with variational recurrent neural networks (VRNN) to estimate content popularity. Through multi-agent deep reinforcement learning, the cooperative edge caching algorithm substantially improved the benefits of edge caches in ultra-dense next-generation networks. In [30], Jiang et al. employed multi-agent reinforcement learning to design content caching strategies in mobile device-to-device (D2D) networks, addressing the challenge of content caching without prior knowledge of popularity distribution. The proposed approach, using Q-learning, maximized caching rewards by treating users as agents and caching decisions as actions. In [31], Wu et al. proposed the integration of cooperative coded caching using maximum distance separable (MDS) codes in small cell networks to alleviate traffic loads. To adapt to dynamic and unknown content popularity, a multi-agent deep reinforcement learning framework was introduced to intelligently update cached content, with a focus on minimizing long-term expected front-haul traffic loads, resulting in improved caching efficiency. In [32], Zhong et al. investigated cache placement by using the collaborative MADRL algorithm in ultra-dense small cell networks (SCNs) to maximize cache hit rates in centralized and decentralized scenario settings. In [33], Jiang et al. explored edge caching for the internet of vehicles (IoVs) using distributed MARL. They introduced a hierarchical caching architecture to minimize content delivery overhead, extended single-agent reinforcement learning to a multi-agent system, and achieved superior performance in comparison to other caching strategies through simulation results. In [37], Song et al. addressed the content caching and sharing problem in cooperative base stations while considering unknown content popularity distribution and content retrieval costs. This paper formulated the problem as a multi-armed bandit (MAB) learning problem, jointly optimizing caching and sharing strategies. Two algorithms, a centralized approach using semidefinite relaxation and a decentralized approach based on the alternating direction method of multipliers (ADMM), were proposed to solve the problem efficiently. In [38], Cai et al. proposed a cross-tier cooperative caching architecture for all content that allows distributed cache nodes to cooperate, using a MADRL approach to model the decision process between heterogeneous cache nodes. In [39], Wang et al. introduced an intelligent edge caching framework named MacoCache, which incorporated MADRL to enhance edge caching performance in distributed environments. By employing the advanced actor-critic method and integrating long short-term memory (LSTM) for time series dynamics, MacoCache successfully reduced latency by an average of 21% and costs by 26% compared to existing learning-based caching solutions. In [40], Chen et al. utilized a refined MADRL tailored for the 5G network structure to improve scalable video coding (SVC)-based caching and video quality service strategies. The introduced value-decomposed dimensional network (VDDN) algorithm and a dimension decomposition method in the dueling deep Q-network addressed the challenges of large action spaces and computational efficiency. In [41], Zhang et al. introduced an edge caching model for IoV content distribution, allowing vehicles to collaboratively cache content and account for varying content popularity and channel conditions. An autonomous decision-making approach, an edge caching approach for the IoV based on multi-agent deep reinforcement learning, was proposed where vehicles act as agents to optimize caching decisions. However, most previous studies have primarily used MADRL to tackle issues in collaborative content caching. It is important to point out that these studies often did not address the more complex challenges associated with predicting what contents will be popular in the future. Specifically, they did not focus on the intricacies of using machine learning to accurately forecast content popularity, which is crucial for effective content management.
@@
Recently, there have been several works on designing content caching schemes based on MADRL and FL. In [42], Chang et al. introduced a federated multi-agent reinforcement Learning (FMRL) method for optimizing edge caching. Using a Markov decision process (MDP) and a specialized regression model, it predicted content popularity, improved server rewards, and optimized cache decision-making. An enhanced FL technique aggregated local models, improving training and preserving user privacy. In [43], Prathiba et al. introduced the federated Learning and edge cache-assisted Cybertwin framework in response to rising autonomous vehicle (AV) technology and mobile traffic. This framework integrated the cybertwin concept to merge physical and digital systems. It employed the federated MADRL algorithm for optimized learning, considering various factors. Additionally, the federated reinforcement learning-based edge caching algorithm was used for efficient caching in 6G vehicle-to-everything networks. In [44], Sun et al. employed FL to perceive user preferences across various districts without exposing user requests. A content blocks-based RL is utilized for intelligent caching, taking into account the receiving scenarios of AVs. By integrating FL with MARL, this paper achieved distributed training of local models, which not only reduced the state-action space dimension but also circumvented local optimal solutions. However, the previous studies did not take into account the personalization needs of individual UE in FL training. Additionally, they overlooked the fact that different SBSs may have various popular contents, which can significantly impact caching strategies.
In conclusion, previous research has not adequately addressed two key aspects in designing content caching schemes. Firstly, FL approaches have not considered the personalized and elastic requirements of each UE, which is essential for optimizing user-specific performance. Secondly, MADRL methods have largely ignored the importance of predicting popular contents before making collaborative caching decisions. These gaps in the literature have motivated our research, as we aim to develop personalized models for UEs while also focusing on proactive collaborative caching strategies based on predicted content popularity.
@@
$\bm{a^{t}}$ the global action at time slot $t$. $a_{b}^{t}$ the action of SBS $b$ at time slot $t$.
$B$ the number of SBSs. $b_{d}$ the bias vector of the decoder network.
$b_{e}$ the bias vector of the encoder network. $C$ the cache capacity of each SBS.
@@
$c_{b}^{t}$ the current contents cached by SBS $b$ at time slot $t$. $d$ the size of entire data of all UEs within the coverage of SBS.
@@
$d(\cdot)$ the non-linear and logically-activated function of the discriminator network. $d_{i}$ the size of the training data of each UE.
$D$ the replay buffer. $D(z)$ the confidence of $z$.
@@
$e$ the number of iteration of the local AAE training for UE. $E^{\prime}$ the number of episodes in the testing phase of MADRL algorithm.
@@
$F_{p}$ the number of predicted popular contents. $J(\bm{\pi^{t}})$ the expected long-term discounted reward at time slot $t$.
$|L|$ the number of layers of the AAE model. $M$ the number of transition tuples in a mini-batch.
@@
$n_{i,k}^{r}$ local training data set of UE $i$ in the iteration $k$ of round $k$. $p(\cdot)$ the non-linear and logically-activated function of the decoder network.
$P(z)$ the predefined distribution. $p_{b}$ the predicted popular contents of SBS $b$ coverage.
@@
$q(\cdot)$ the non-linear and logically-activated function of the encoder network. $R^{t}$ the global reward at the time slot $t$.
@@
$R_{b}^{t}$ the local reward of SBS $b$ at time slot $t$. $R_{max}$ the number of rounds to train the AAE model in the elastic FL.
@@
$r_{b}^{r}(t)$ the number of contents that SBS $b$ fetches from the local cache at time slot $t$. $r_{b,c}^{r}(t)$ the number of contents that SBS $b$ fetches from CS at time slot $t$.
@@
$r_{b,e}^{r}(t)$ the contents replaced in SBS $b$ during time slot $t$. $r_{b,n}^{r}(t)$ the number of contents that SBS $b$ fetches from adjacent SBSs at time slot $t$.
$s_{b}^{t}$ the local state of SBS $b$ at time slot $t$. $\bm{s^{t}}$ the global state at time slot $t$.
@@
$T$ the number of time slots of the MADRL algorithm. $X_{i}^{r}$ the rating matrix of UE $i$ in round $r$.
@@
$\tilde{X}_{i}^{r}$ the reconstructed rating matrix of UE $i$ in round $r$. $\alpha$ the cost of fetching the contents from local SBS $b$.
@@
$\alpha_{i}$ the elastic parameter of UE $i$. $\beta$ the cost that the UE within the coverage of local SBS $b$ fetch the contents from adjacent SBSs.
$\chi$ the cost of fetching the contents from CS. $\delta$ the cost of replacing a content.
$\gamma$ the discount factor. $\eta$ the fixed learning rate.
@@
$\theta_{b}$ the parameters of the actor network of agent $b$. $\theta_{b}^{\prime}$ the parameters of the target critic network of agent $b$.
$\varphi$ the parameters of the global critic network. $\varphi^{\prime}$ the parameters of the target global critic network.
@@
$\tau$ the constant is used to update the target networks. $\omega^{r}$ the global AAE model in round $r$.
$\omega_{e}^{r}$ the global encoder network in round $r$. $\omega_{d}^{r}$ the global decoder network in round $r$.
@@
$\omega_{d^{\prime}}^{r}$ the global discriminator network in round $r$. $\omega_{i}^{r}$ the local model of UE $i$ in round $r$.